OpenStudy (anonymous):

The ages of students in a school are normally distributed with a mean of 15 years and a standard deviation of 2 years. Approximately what percent of the students are between 14 and 18 years old? 24.17% 62.47% 30.85% 93.32%

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (janu16):

Well, we have : mean = 15 years SD = 2 years let be F the distribution of the N(0,1) is the z-table, then P( 14 < Age < 18) = P( - 0.5 < (Age - mean)/SD < 1.5) = F(1.5) - F( -0.5) = .93319 - .30854 -----------------------> P( 14 < Age < 18) = .62465 second answer

OpenStudy (cbarredo1):

B. 62.47

OpenStudy (anonymous):

Wait wait wait, Can someone tell me exactly how to do this as i have 2 more questions like this one and i should probably figure out how to do them properly

OpenStudy (anonymous):

Pizza delivery times at Pizza Time are normally distributed with a mean time of 27 minutes and a standard deviation of 3 minutes. Approximately what percent of pizzas are delivered between 24 and 30 minutes? 32% 68% 95% 99.7%

OpenStudy (anonymous):

this is the next one i have to do.